Fermi National Laboratory


Accelerator Update for October 28, 2005 - October 31, 2005

Friday October 28
The day shift began with Operations monitoring store 4473 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with NuMI, MiniBooNE, and the Meson MCenter experiments taking beam.

At 8:31 AM, in different combinations, a group of three Main Injector (MI) vacuum pumps (IP7100, 7091, 7092) tripped off for the first of a bunch of times over the next seven hours. They reset okay each time.

At 9:10 AM, a crate power supply at the AP0 transfer line service building tripped off and caused a network link break. Operators discovered that a breaker was inadvertently switched off; this kept beam off until 9:22 AM.

A Tevatron Electron Lens (TEL) expert began an end-of-store study at 10:02 AM. The study ended at 12:17 PM.

Operations terminated store 4473 at 12:28 PM.

Cryo system technicians began maintenance work on the Tevatron (TeV) sector E2 system at 12:59 PM.

At 1:35 PM, the TeV suffered a sector F2 & F3 quench. This was due to a proton pilot beam being injected at the wrong place.

Operations made the CDF and D0 collision halls ready for an access at 1:49 PM.

The TeV sector A3 cryo work ended at 3:30 PM.

Both CDF and D0 were out of access by 3:45 PM.

Operations put the TeV through a dry squeeze at 4:53 PM.

Operations put the TeV into shot setup at 5:28 PM.

Operations established Recycler only store 4475 at 8:45 PM with an initial luminosity of 138.35E30.

Operations began preparing the Recycler to stash antiprotons at 8:45 PM. The stash ended and Pbar resumed stacking around 9:44 PM.

At 11:44 PM, a Switchyard (SY) power supply (S:Q100) tripped off and wouldn't reset. Operators called in a power supply expert. This failure shut off all Meson beam.

Saturday October 29
The midnight shift began with Operations monitoring store 4475 and the stash, with Pbar stacking, with Meson beam off due to a SY power supply problem, and with MiniBooNE and NuMI taking beam.

The power supply expert found a fuse fault indication in the SY power supply, but the fuse was fine. He reset the indicator at 12:39 AM.

The Recycler began stashing at 1:44 AM. The stash ended at 2:23 AM. Pbar and MCenter resumed taking beam by 2:36 AM.

The Recycler began preparing to stash at 7:03 AM. The stash ended; Pbar and MCenter resumed taking beam by 7:48 AM.

At 9:26 AM, a Pbar Accumulator RF station (ARF1) tripped off and wouldn't reset. An operator contacted an expert and per instructions cleaned the air filters, which allowed the station to reset. Pbar resumed stacking at 11:04 AM.

The Recycler began preparing to stash at 1:53 PM. The stash ended; Pbar and MCenter resumed normal operations at 2:43 PM.

At 7:27 PM, Operations reported that MIRF16 was repeatedly tripping off. Operators contacted a RF expert. The station was reset by 8:06 PM.

TeV experts began an end-of-store study at 10:13 PM.

Operations terminated store 4475 at 11:42 PM. An LCW expert used this time to replace the TeV sector C2 LCW filters.

Operations put the TeV into shot setup at 11:51 PM.

Sunday October 30
The midnight shift began with Operations monitoring the stash, with TeV in shot setup, with Pbar and MCenter off due to shot setup, and with MiniBooNE and NuMI taking beam.

At 1:08 AM, Operations reported that all the sequencers were faulting on state device problems. This occurred at the time change.

Operations managed to successfully reboot the state processes at 1:29 AM.

Operations established store 4476 at 1:52 AM with an initial luminosity of 148.5E30. Pbar resumed stacking.

MCenter resumed taking beam at 2:24 AM. Operations also reported ACNET and Datalogging problems due to the time change.

The Recycler began preparing to stash at 3:25 AM. The stash ended; Pbar and MCenter resumed normal operations at 4:14 AM.

The Recycler began preparing to stash at 8:06 AM. The stash ended; Pbar and MCenter resumed normal operations by 9:12 AM.

The Recycler began preparing to stash at 1:30 PM. At 1:52 PM, two Recycler vacuum valves closed due to bad vacuum. With advice from an expert, operators reset the vacuum pumps, the vacuum improved, and the stash continued at 2:01 PM. The stash ended; Pbar and MCenter resumed normal operations by 2:28 PM.

The Recycler began preparing to stash at 5:44 PM. The stash ended; Pbar and MCenter resumed normal operations by 6:24 PM.

The Recycler began preparing to stash at 9:25PM. The stash ended; Pbar and MCenter resumed normal operations by 10:21 PM.

Monday October 31
The midnight shift began with Operations monitoring store 4476 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing to stash at 4:56 AM. The stash ended; Pbar and MCenter resumed normal operations by 5:57 AM.

Operations terminated store 4476 at 7:03 AM.

Operations put the TeV into shot setup at 7:33 AM.

The Plans for Monday
The plans for today are to stack, stash, and store, and to allow Booster and MI experts to conduct studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/31/2005   email Fermilab